# An Incremental Unified Framework for Small Defect Inspection

### 14 Dec 2023 v1

@@ Abstract
Artificial Intelligence (AI)-driven defect inspection is pivotal in industrial manufacturing. Yet, many methods, tailored to specific pipelines, grapple with diverse product portfolios and evolving processes. Addressing this, we present the Incremental Unified Framework (IUF) that can reduce the feature conflict problem when continuously integrating new objects in the pipeline, making it advantageous in object-incremental learning scenarios. Employing a state-of-the-art transformer, we introduce Object-Aware Self-Attention (OASA) to delineate distinct semantic boundaries. Semantic Compression Loss (SCL) is integrated to optimize non-primary semantic space, enhancing network adaptability for novel objects. Additionally, we prioritize retaining the features of established objects during weight updates. Demonstrating prowess in both image and pixel-level defect inspection, our approach achieves state-of-the-art performance, proving indispensable for dynamic and scalable industrial inspections.
@@ 1 Introduction
While most of the current inspection systems [10, 18, 19, 9] are specifically designed for particular industrial products (One-Model-One-Object, cf. Fig. 1 (A) (B)), the ever-changing dynamics of real-world product variants present two salient challenges. Firstly, there is a pressing need for a system capable of detecting multiple objects. Secondly, the adaptability of these systems to the frequently adjusted production schedules [29] is a lingering concern. Relying solely on acquiring new inspection equipment is neither cost-effective nor efficient due to increased deployment time.
@@
Figure 1: Different framework in small defect inspection. (A) shows the most common One-Model-One-Object pattern, which trains a separate model for each of the different objects. (B), based on (A), the types of defects are incrementing, which improves the generalization performance in detecting different defects. (C) shows a unified model for multi-objects. (D) use a memory bank to incrementally record features for all objects for distinguishing. (E) is our Incremental Unified framework, and it combines the advantages of both (C) and (D).
@@
Figure 2: Semantic space in our methodology (B-D). In (A), all objects in the original semantic space are tightly coupled, leading to “catastrophic forgetting", i.e., the learning of new objects will result in forgetting previously learned objects. Our method firstly builds a semantic boundary of each object (B), then compact the non-primary semantic space (C), and finally suppress semantic updates of the primary object’s features (D).
@@
Currently, only limited research addresses these challenges. One-Model-N-Object approaches like UniAD [34] and OmiAD [37] proposees a unified network suitable for a broad spectrum of objects. However, its capability is restricted when encountering dynamic or unfamiliar objects. On the other hand, CAD [18] suggests embedding features of multiple objects into a memory bank, with a binary classifier discerning defects (Fig. 1 (D)). Yet, the differences in object distributions can jeopardize memory bank features, escalating storage demands and impeding pixel-level performance evaluation.
@@
In an attempt to surmount the confines of prior research, we introduce an Incremental Unified Framework (IUF) which integrates the multi-object unified model with object-incremental learning, as depicted in Fig. 1 (E). IUF capitalizes on the unified model’s prowess, enabling pixel-precise defect inspection across diverse objects without necessitating an embedded feature storage in a memory bank.
@@
In this framework, one significant challenge is “catastrophic forgetting", which hampers network efficacy in the Incremental Unified framework due to semantic feature conflicts in the reconstruction network (Fig. 2 (A)). We propose the Object-Aware Self-Attention (OASA), leveraging object category features to segregate the semantic spaces of different objects (Fig.2 (B)). These features serve as semantic constraints within the encoding process of the reconstruction network, establishing distinct feature boundaries in the semantic hyperplane, thereby alleviating feature coupling. Then, we introduce Semantic Compression Loss (SCL) to condense non-essential feature values, concentrating the feature space on principal components (Fig.2 (C)). This creates more semantic space for optimizing unseen objects and minimizing potential forgetting issues. Finally, when learning novel objects, we propose a new updating strategy to retain features of prior objects and reduce interference from new objects in the prevailing feature space (Fig.2 (D)). Our framework offers a versatile solution for small defect inspection, adeptly navigating the challenges posed by feature space conflicts, and achieves state-of-the-art (SOTA) performance at both image and pixel levels.
@@ 2.1 Small Defect Inspection
Another more efficient method is the reconstruction-based approach that operates on an alternative hypothesis. Such methods posit that models trained on normal samples excel in reproducing normal regions while struggling with anomalous regions [4, 8, 34, 37]. Therefore, many approaches aim to discover an optimal network architecture for effective sample reconstruction, including Autoencoder (AE) [4], Variational Autoencoder (VAE) [19], Generative Adversarial Networks (GANs) [1], Transformer [34], etc. Furthermore, beyond image-level reconstruction, certain methods endeavor to reconstruct features at a more granular level, e.g., InTra [24] and UniAD [34]. Until recently, the majority of prior approaches adhered to the One-Model-One-Object pattern as shown in Fig. 1 (A). However, recent advancements such as UniAD [34] and OmniAL [37] have extended the reconstruction paradigm to the One-Model-N-Objects pattern (depicted in Fig. 1 (C)). This progress curtails memory usage compared to the one-model-one-object approach, propelling the entire problem into a more general framework.
@@ 2.2 Incremental Learning for Defect Inspection
Incremental learning (also known as continuous learning, lifelong learning) aims to enable AI systems to continuously acquire, update, accumulate, and utilize knowledge in response to external changes [30]. There have been some previous studies combining incremental learning and small defect inspection, but these studies mainly focused on augmenting the defect types of an Object under one specific scenarios [33, 5, 28, 27, 7]. The purpose of these methods is to improve the performance of the model. Most recently, CAD [18] initially proposed an unsupervised feature embedding-based method for continuous learning of task sequences, but there are several problems.
@@ 3 Problem Formulation
To conduct reconstruction-based small defect inspection [34], we need an autoencoder for reconstructing normal features in all objects, as
@@
Figure 3: Problems in our framework. (A-1) and (A-2) demonstrate the performance of image-level and pixel-level models under the current framework, where catastrophic forgetting significantly occurs. (B) demonstrates the reason for catastrophic forgetting. When training the current step, the training model overwrites the previous semantic patterns, causing severe feature conflicts in the reconstructed network.
@@
Under Eq. (2), small defect inspection is also notably susceptible to catastrophic forgetting, as shown in Fig. 3 (A-1) and (A-2). Unlike other incremental learning tasks, small defect inspection requires the reconstruction of normal features, necessitating the decoding of spatial information from multiple channels. Without the rehearsal of the old category, the features of the new category will modify and occupy the original feature channel, resulting in semantic conflicts in space, as shown in Fig. 3.
Intuitively, our objective is to reorganize distinct categories of features into their respective channels, all the while avoiding any potential semantic conflicts. Nevertheless, within our framework, the task of establishing meaningful semantic relationships between these diverse categories presents a considerable challenge.
@@ 4.1 Identifying Semantics Boundary
where $\cdot$ is Hadamard product, $d_{k}$ is the dimension of the key or query, and $T$ is the matrix transpose operator. $Q\in\mathbb{R}^{T\times C\times H\times W}$ is the query in $T$ key layers of the transformer-based network. $K$ and $V$ is the key and value of this network. By inserting $C_{o_{n}}$ to $Q$, the reconstruction network can identify semantic boundaries explicitly.
@@ 4.2 Compacting Semantic Space
Specifically, the latent features $M_{n}\in\mathbb{R}^{B\times C\times H\times W}$ is firstly compressed by aggregating spatial features as in Eq. (5),
$\hat{M}_{n}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}M_{n}(b,c,h,w),$ (5)
where $\hat{M}_{n}\in\mathbb{R}^{B\times C}$ represents the semantic information matrix of latent features in each batch. As is common in the field [11, 23], we hypothesize that different objects of semantic features are distributed on distinct channels, while spatial information is relatively irrelevant. Then, we perform SVD on the first two dimensions $B$ and $C$ of $\hat{M}_{n}$. The representation formula is as Eq. (6),
$\hat{M}_{n}$ $=U_{n}S_{n}V_{n}^{T}$ (6)
$=U_{n}\begin{bmatrix}\sigma_{1}&0&\cdots&0\\ 0&\sigma_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sigma_{k}\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ \end{bmatrix}V_{n}^{T},$
where three matrices obtained after SVD, i.e., $U_{n}$, $S_{n}$, and $V_{n}^{T}$. Among them, $U\in\mathbb{R}^{B\times B}$ is an orthogonal matrix for batch eigenspace, $S_{n}=\mathrm{diag}(\sigma_{1},\sigma_{2},...,\sigma_{k})\in\mathbb{R}^{B\times C}$ is a diagonal matrix of eigenvalues, and $V^{T}\in\mathbb{R}^{C\times C}$ is an orthogonal matrix of channel eigenspace.
$S_{n}$ reflects the importance of semantic information for $M_{c}$ on each channel dimension. Larger eigenvalues correspond to more important semantic information, while smaller eigenvalues correspond to relatively non-primary semantic information. Based on this fact, we can compress the features of each object by reducing the non-primary semantic information. Therefore, we construct the Semantic Compression Loss (SCL) as in Eq. (7),
$L_{sc}=\sum_{i=t}^{k}\sigma_{i},$ (7)
where $t\in[1,k]$ is a hyperparameter, which is a scale factor that reflects the degree of compression of the semantic feature space. To construct the total loss, we use $L=\lambda_{0}L_{1}+\lambda_{1}L_{CE}+\lambda_{2}L_{sc}$, $\lambda_{0}$, $\lambda_{1}$ and $\lambda_{2}$ is hyperparameter for total loss. This loss continuously compacts the non-primary feature space while achieving the final optimization goal.
@@ 4.3 Reinforcing Primary Semantic Memory
$\nabla\theta_{j}$ $=-\alpha\frac{\partial L(\theta)}{\partial\theta_{j}}$ (8)
@@
Retaining Prior Semantic Information To maintain the semantics of the old objects, it is essential to continuously consolidate the existing weights in old objects during the incremental process. Therefore, we continuously copy the old weight in gradient descent in the next step, as in Eq. (9),
@@
Based on this motivation, we consider that the importance of the semantic space can be represented by $V_{n}^{T}$ (Eq. (6)). Thus, we project the updating weight, $\nabla\theta_{j}$, to the corresponding channel space in old objects, as in Eq. (10),
@@
where $\nabla\Theta_{j}$ is updating weight in channel space. $V_{old}^{T}$ is channel eigenspace in old objects. To constrain the updating in primary semantic space in old objects, we empirically use a $\log$ function to constrain the updating of different channels, $c\in\mathbb{R}^{1\times C}$, as Eq. (11),
$\Omega(k,c)=k*\log(c)$ (11)
where $\Omega(\cdot)$ is constraining function. By acting on this function to $\nabla\Theta_{j}$ and then projecting it back to the original space, we finally get the updating weight in the original space, as in Eq. (12),
@@
Table 1: Qualitative evaluation. (A) and (B) show the performance at the image and pixel-level respectively. “N/A" is Not Available, CAD [18]-based method cannot locate defects in pixel-level. Red represents the best ACC.
Input UniAD UniAD + EWC Ours Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative Evaluation in $\mathbf{10-1\times 5\ (5\ Steps)}$. Compared with other reconstruction-based baselines, our method can significantly reduce semantic feature conflict and locate defects more accurately.
5 Experiment
@@ 5.1 Experimental Setup
Dataset We choose MVTec-AD [3] as our dataset, which is a real-world industrial small defect inspection dataset and widely used in this task [18, 19, 34]. MVTec-AD contains 15 types of objects or textures, so it can be available for our Incremental Unified Framework.
Baselines We select some SOTA methods in small defect inspection as the baseline, including PaDiM [10], DRAEM [35], PANDA [25], CutPaste [17], and UniAD [34]. Notably, some baselines integrate with the incremental learning protocol, CAD [18]. In addition, we integrate several available SOTA baselines in incremental learning with UniAD, including EWC [16], SI [36], MAS [2] and LVT [31].
Protocol According to our framework in Fig. 1 (E), our reconstruction model incrementally learns new objects. When it learns at the current step, the data from the previous steps are not available. Therefore, based on the practical requirements of industrial defect inspection, we set up our experiments in both single-step and multi-step settings. In single-step, we conduct the following experiments: $\mathbf{14-1\ (1\ Step)}$, $\mathbf{10-5\ (1\ Step)}$. In multi-step, we conduct the following experiments: $\mathbf{3\times 5\ (5\ Steps)}$, $\mathbf{10-1\times 5\ (5\ Steps)}$.
Evaluation Metrics Currently, there are two metrics for continual learning: average accuracy (ACC) [21] and forgetting measure (FM) [6]. However, for individual tasks, we usually use Pixel-level AUROC, $A^{pixel}$, and Image-level AUROC, $A^{image}$, to characterize the detection accuracy. In this task, we combine these two goals and define four metrics, as in Eq. (14) (ACC) and Eq. (15) (FM),
$\frac{1}{N}\sum_{i=1}^{N}A^{pixel}_{N,i}\ \textbf{and}\ \frac{1}{N}\sum_{i=1}^{N}A^{image}_{N,i},$ (14)
$\frac{1}{N-1}\sum_{i=1}^{N-1}\max_{b\in\{1,\cdots,N-1\}}(A^{pixel}_{b,i}-A^{pixel}_{N,i})$ (15)
  and $\frac{1}{N-1}\sum_{i=1}^{N-1}\max_{b\in\{1,\cdots,N-1\}}(A^{image}_{b,i}-A^{image}_{N,i}).$

### 24 Jan 2024 v2

@@ Abstract
Artificial Intelligence (AI)-driven defect inspection is pivotal in industrial manufacturing. Yet, many methods, tailored to specific pipelines, grapple with diverse product portfolios and evolving processes. Addressing this, we present the Incremental Unified Framework (IUF), which can reduce the feature conflict problem when continuously integrating new objects in the pipeline, making it advantageous in object-incremental learning scenarios. Employing a state-of-the-art transformer, we introduce Object-Aware Self-Attention (OASA) to delineate distinct semantic boundaries. Semantic Compression Loss (SCL) is integrated to optimize non-primary semantic space, enhancing network adaptability for novel objects. Additionally, we prioritize retaining the features of established objects during weight updates. Demonstrating prowess in both image and pixel-level defect inspection, our approach achieves state-of-the-art performance, proving indispensable for dynamic and scalable industrial inspections.
@@ 1 Introduction
While most of the current inspection systems [10, 18, 19, 9, 25, 33, 5, 28, 27, 7] are specifically designed for particular industrial products (One-Model-One-Object, cf. Fig. 1 (A) (B)), the ever-changing dynamics of real-world product variants present two salient challenges. Firstly, there is a pressing need for a system capable of detecting multiple objects. Secondly, the adaptability of these systems to the frequently adjusted production schedules [29] is a lingering concern. Relying solely on acquiring new inspection equipment is neither cost-effective nor efficient due to increased deployment time.
@@
Figure 1: Different framework in small defect inspection. (A) shows the most common One-Model-One-Object pattern [10, 18, 19, 9, 25], which trains a separate model for each of the different objects. (B), based on (A), the types of defects are incrementing [33, 5, 28, 27, 7], which improves the generalization performance in detecting different defects. (C) shows a unified model [34, 37] for multi-objects. (D) use a memory bank to incrementally record features for all objects for distinguishing [18]. (E) is our Incremental Unified framework, and it combines the advantages of both (C) and (D).
@@
Figure 2: Semantic space in our methodology (B-D). In (A), all objects in the original semantic space are tightly coupled, leading to “catastrophic forgetting", i.e., the learning of new objects will result in forgetting previously learned objects. Our method firstly builds a semantic boundary of each object (B), then compacts the non-primary semantic space (C), and finally suppresses semantic updates of the primary object’s features (D).
@@
Currently, only limited research addresses these challenges. One-Model-N-Object approaches like UniAD [34] and OmiAD [37] propose a unified network suitable for a broad spectrum of objects. However, its capability is restricted when encountering dynamic or unfamiliar objects. On the other hand, CAD [18] suggests embedding features of multiple objects into a memory bank, with a binary classifier discerning defects (Fig. 1 (D)). Yet, the differences in object distributions can jeopardize memory bank features, escalating storage demands and impeding pixel-level performance evaluation.
@@
In an attempt to surmount the confines of prior research, we introduce an Incremental Unified Framework (IUF), which integrates the multi-object unified model with object-incremental learning, as depicted in Fig. 1 (E). IUF capitalizes on the unified model’s prowess, enabling pixel-precise defect inspection across diverse objects without necessitating an embedded feature storage in a memory bank.
@@
In this framework, one significant challenge is “catastrophic forgetting", which hampers network efficacy in the Incremental Unified framework due to semantic feature conflicts in the reconstruction network (Fig. 2 (A)). We propose Object-Aware Self-Attention (OASA), leveraging object category features to segregate the semantic spaces of different objects (Fig.2 (B)). These features serve as semantic constraints within the encoding process of the reconstruction network, establishing distinct feature boundaries in the semantic hyperplane, thereby alleviating feature coupling. Then, we introduce Semantic Compression Loss (SCL) to condense non-essential feature values, concentrating the feature space on principal components (Fig.2 (C)). This creates more semantic space for optimizing unseen objects and minimizing potential forgetting issues. Finally, when learning novel objects, we propose a new updating strategy to retain features of prior objects and reduce interference from new objects in the prevailing feature space (Fig.2 (D)). Our framework offers a versatile solution for small defect inspection, adeptly navigating the challenges posed by feature space conflicts, and achieves state-of-the-art (SOTA) performance at both image and pixel levels.
@@ 2.1 Small Defect Inspection
Another more efficient method is the reconstruction-based approach that operates on an alternative hypothesis. Such methods posit that models trained on normal samples excel in reproducing normal regions while struggling with anomalous regions [4, 8, 34, 37]. Therefore, many approaches aim to discover an optimal network architecture for effective sample reconstruction, including Autoencoder (AE) [4], Variational Autoencoder (VAE) [19], Generative Adversarial Networks (GANs) [1], Transformer [34], etc. Furthermore, beyond image-level reconstruction, certain methods endeavor to reconstruct features at a more granular level, e.g., InTra [24] and UniAD [34]. Until recently, the majority of prior approaches adhered to the One-Model-One-Object pattern, as shown in Fig. 1 (A). However, recent advancements such as UniAD [34] and OmniAL [37] have extended the reconstruction paradigm to the One-Model-N-Objects pattern (depicted in Fig. 1 (C)). This progress curtails memory usage compared to the one-model-one-object approach, propelling the entire problem into a more general framework.
@@ 2.2 Incremental Learning for Defect Inspection
Incremental learning (also known as continuous learning, lifelong learning) aims to enable AI systems to continuously acquire, update, accumulate, and utilize knowledge in response to external changes [30]. There have been some previous studies combining incremental learning and small defect inspection, but these studies mainly focused on augmenting the defect types of one object under one specific scenario [33, 5, 28, 27, 7]. The purpose of these methods is to improve the performance of the model. Most recently, CAD [18] initially proposed an unsupervised feature embedding-based method for continuous learning of task sequences, but there are several problems.
@@ 3 Problem Formulation
For small defect inspection by reconstruction methods [34], an autoencoder is required to reconstruct normal features in all objects, as Eq. 1,
@@
[…]
@@
In such task pipelines, we notice that small defect inspection is also notably susceptible to catastrophic forgetting, as shown in Fig. 3 (A-1) and (A-2). Unlike other incremental learning tasks, small defect inspection requires the reconstruction of normal features, necessitating the decoding of semantic information from multiple channels. Without the rehearsal of the old category, the features of the new category will modify and occupy the original feature channel, resulting in semantic conflicts in space, as shown in Fig. 3.
Intuitively, our objective is to reorganize distinct categories of features into their respective channels, while avoiding any potential semantic conflicts. Nevertheless, within our framework, the task of establishing meaningful semantic relationships between these diverse categories presents a considerable challenge.
@@ 4.1 Identifying Semantics Boundary
where $\cdot$ is Hadamard product, $d_{k}$ is the dimension of the key or query, and $T$ is the matrix transpose operator. $Q\in\mathbb{R}^{T\times C\times H\times W}$ is the query in $T$ key layers of the transformer-based network. $K$ and $V$ are the keys and values of this network, respectively. By inserting $C_{o_{n}}$ to $Q$, the reconstruction network can identify semantic boundaries explicitly.
@@ 4.2 Compacting Semantic Space
Specifically, the latent features $M\in\mathbb{R}^{B\times C\times H\times W}$ is firstly compressed by aggregating spatial features as in Eq. (5),
$\hat{M}=\frac{1}{H\times W}\sum_{h=1}^{H}\sum_{w=1}^{W}M_{n}(b,c,h,w),$ (5)
where $\hat{M}\in\mathbb{R}^{B\times C}$ represents the semantic information matrix of latent features in each batch. As is common in the field [11, 23], we hypothesize that different objects of semantic features are distributed on distinct channels, while spatial information is relatively irrelevant. Then, we perform SVD on the first two dimensions $B$ and $C$ of $\hat{M}$. The representation formula is as Eq. (6),
$\hat{M}$ $=USV^{T}$ (6)
=[$=\underset{\mathbf{Batch\ Space}}{\underbrace{\left[\begin{matrix}u_{11}&u_{12}&u_{13}&\cdots&u_{1B}\\ u_{21}&u_{22}&u_{23}&\cdots&u_{2B}\\ u_{31}&u_{32}&u_{33}&\cdots&u_{3B}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ u_{B1}&u_{B2}&u_{B3}&\cdots&u_{BB}\\ \end{matrix}\right]}}\left[\begin{matrix}\sigma_{1}&0&\cdots&0\\ 0&\sigma_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\sigma_{C}\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ \end{matrix}\right]\underset{\mathbf{Channel\ Space}}{\underbrace{\left(\left[\begin{matrix}v_{11}&v_{12}&\cdots&v_{1C}\\ v_{21}&v_{22}&\cdots&v_{2C}\\ \vdots&\vdots&\ddots&\vdots\\ v_{C1}&v_{C2}&\cdots&v_{CC}\\ \end{matrix}\right]\right)^{T}}},$
where three matrices obtained after SVD, i.e., $U$, $S$, and $V^{T}$. Among them, $U\in\mathbb{R}^{B\times B}$ is an orthogonal matrix for batch eigenspace, $S=\mathrm{diag}(\sigma_{1},\sigma_{2},...,\sigma_{C})\in\mathbb{R}^{B\times C}$ is a diagonal matrix of eigenvalues, and $V^{T}\in\mathbb{R}^{C\times C}$ is an orthogonal matrix of channel eigenspace.
$S_{n}$ reflects the importance of semantic information for $M$ on each channel dimension. Larger eigenvalues correspond to more important semantic information, while smaller eigenvalues correspond to relatively non-primary semantic information. Based on this fact, we can compress the features of each object by reducing the non-primary semantic information. Therefore, we construct the Semantic Compression Loss (SCL) as in Eq. (7) and Fig. 5,
$L_{sc}=\sum_{i=t}^{C}\sigma_{i},$ (7)
where $t\in[1,C]$ is a hyperparameter, which is a scale factor that reflects the degree of compression of the semantic feature space. To construct the total loss, we use $L=\lambda_{0}L_{1}+\lambda_{1}L_{CE}+\lambda_{2}L_{sc}$, where $\lambda_{0}=1$, $\lambda_{1}=0.5$ and $\lambda_{2}$ is hyperparameter for total loss. $\lambda_{2}$ is set between $1-10$ based on the experiment. This loss continuously compacts the non-primary feature space while achieving the final optimization goal.
@@ 4.3 Reinforcing Primary Semantic Memory
$\nabla\theta_{j}$ $=-\alpha\frac{\partial L(\theta)}{\partial\theta_{j}},$ (8)
@@
Retaining Prior Semantic Information To maintain the semantics of the old objects, it is essential to continuously consolidate the existing weights in old objects during the incremental process. Therefore, we constantly copy the old weight in gradient descent in the next step, as in Eq. (9),
@@
Based on this motivation, we consider that the importance of the semantic space can be represented by $V^{T}$ (Eq. (6)). Thus, we project the updating weight, $\nabla\theta_{j}$, to the corresponding channel space in old objects, as in Eq. (10),
@@
where $\nabla\Theta_{j}\in\mathbb{R}^{C\times W}$ is updating weight in channel space. $V_{old}^{T}$ is channel eigenspace in previous steps.
To constrain the updating in primary semantic space in previous steps, we empirically use a $\log$ function to constrain the updating of different channels, $c\in\mathbb{R}^{1\times C}$, as Eq. (11),
$\Omega(k,c)=k\times\log(c),$ (11)
where $\Omega(\cdot)$ is constraining function. When the channel index $c$ equals $1$, the result of the $\log$ function is 0, indicating that the model does not update in this specific dimension. This is because the model is in the most crucial semantic space for the previous objects. As the channel index $c$ increases, this channel becomes more likely to trigger model updates, and these dimensions become less important for representing previous objects.
By acting on this function to $\nabla\Theta_{j}$ and then projecting it back to the original space, we finally get the updating weight in the original space, as in Eq. (12) and Fig. 5,
@@ 5 Experiments
(B) Quantitative Performance in VisA [39].
[…]
Table 1: Quantitative evaluation in MvTec AD [3] (A) and VisA [39] (B). “Image / Pixel" shows the image-level and pixel-level performance respectively. “NA" is Not Available, since CAD [18]-based method cannot locate defects in pixel-level. Red and Gray Red represents the best image-level and pixel-level performance respectively.
@@ 5.1 Experimental Setup
Datasets We choose MVTec-AD [3] and VisA [39] as our dataset. MVTec-AD [3] and VisA [39] have 15 and 12 types of objects, respectively. These two employed datasets have well-recognized complex cases for real-world evaluation, and different objects’ feature domains show significantly varying distributions, so both of them can be available for our Incremental Unified Framework.
[…]
Baselines We select some SOTA methods in small defect inspection as the baseline, including PaDiM [10], DRAEM [35], PatchCore [26], PANDA [25], CutPaste [17], and UniAD [34]. Notably, some baselines integrate with the incremental learning protocol, CAD [18]. In addition, we integrate several available SOTA baselines in incremental learning with UniAD, including EWC [16], SI [36], MAS [2] and LVT [31].
Input UniAD [34] UniAD + SI [36] UniAD + MAS [2] UniAD + LVT [31] UniAD + EWC [16] Ours Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative Evaluation. Compared with other baselines, our method can significantly reduce semantic feature conflict and locate defects more accurately.
Evaluation Metrics Currently, there are two metrics for continual learning: average accuracy (ACC) [21] and forgetting measure (FM) [6]. However, for individual tasks, we usually use Pixel-level AUROC, $A^{\mathbf{pix}}$, and Image-level AUROC, $A^{\mathbf{img}}$, to characterize the detection accuracy. In this task, we combine these two goals and define four metrics, as in Eq. (14),
$\small ACC=\left\{\begin{aligned} &\frac{1}{N}\sum_{i=1}^{N-1}{A_{N,i}^{\mathbf{pix}}}\,\,\\ &\frac{1}{N}\sum_{i=1}^{N-1}{A_{N,i}^{\mathbf{img}}}\\ \end{aligned}\right.,FM=\left\{\begin{aligned} &\frac{1}{N-1}\sum_{i=1}^{N-1}{\max_{b\in\{1,\cdots,N-1\}}}\left(A_{b,i}^{\mathbf{pix}}-A_{N,i}^{\mathbf{pix}}\right)\\ &\frac{1}{N-1}\sum_{i=1}^{N-1}{\max_{b\in\{1,\cdots,N-1\}}}\left(A_{b,i}^{\mathbf{img}}-A_{N,i}^{\mathbf{img}}\right)\\ \end{aligned}\right..$ (14)